# Conformal prediction for uncertainties in nucleon-nucleon scattering

### v1

@@ III.1 Pointwise Model
The empirical coverage values closely follow the ideal line, always remaining within the 68% and 95% confidence band. These results validate the robustness and reliability of our method and show that conformal prediction provides guaranteed uncertainty bands with minimal assumptions.
@@ III.2 Gaussian Processes
Here, $l$ is the length scale, which controls how quickly the function can change. In addition, a scaled inverse-chi-squared prior is considered for the $\overline{c}^{2}$, which is a conjugate prior for the variance of the GP. For more details about choosing the priors and hyperparameters in GP, refer to Furnstahl et al. (2015b); Melendez et al. (2017, 2019). Here, we use the same data as applied in the pointwise model Scattering cross section data taken from .
Figure 13: CP intervals for the differential cross section as a function of scattering angle, computed at N2LO and $E=96\ \text{MeV}$ using EKM potential with $R=0.9\ \text{fm}$. The dark and light blue bands represent the $68\%$ and $95\%$ CP intervals, respectively. The black curve corresponds to the Nijmegen PWA result Stoks et al. (1993). The prediction bands are obtained by sampling the chiral EFT expansion coefficients using the GP model and applying CP to construct distribution-free prediction intervals constructed from 10000 samples.
After sampling the total cross-section expansion coefficients at N2LO from the GP model across different ranges of energies, we draw samples for the total cross-section $\sigma_{tot}$ at each energy point. Using these samples, we then apply CP to construct a distribution-free prediction interval. Figure 9(b) illustrates the resulting 68% and 95% CP bands for the total cross section at N2LO order using the EKM potential with $R=0.9\ \text{fm}$. The conformal bands widen gradually with increasing energy, capturing the uncertainty associated with the EFT predictions at higher energy. Looking at Fig. 9, we can compare the CP intervals for the total cross section $\sigma_{tot}$ at the N2LO order obtained using the pointwise model (left panel) and GP model (right panel), across a range of energies. As we can see, both models yield prediction intervals that widen with energy. However, a notable difference is that the GP-based intervals are generally narrower than those from the pointwise model, particularly at higher energies. Additionally, CP bands based on GP are larger than those of the pointwise model for lower energies.
To further explore the versatility of CP, we apply it to the differential cross section $\frac{d\sigma}{d\Omega}$ as a function of scattering angle at a fixed energy of $E=96\ \text{MeV}$. For this, we model the differential cross-section expansion coefficients as random functions using GP, allowing us to generate samples of the total cross sections across the whole angle range. Figure 13 presents the resulting $68\%$ and $95\%$ CP intervals. These results highlight the ability of CP to provide robust, distribution-free uncertainty quantification for a variety of nuclear observables beyond total cross sections. We emphasize that this approach can be generalized to other observables of interest.
@@
Figure 15: 95% and 68% CP intervals for scattering phase shifts in several partial wave channels as a function of energy, generated using the N2LO interactions as described in the main text. The CP bands are constructed from 2500 samples and adapt to the spread of the data, with the black curve showing the Nijmegen partial-wave analysis Stoks et al. (1993).
Figure 16: Comparison of our 95% conformal prediction interval (blue shaded region) with the 95% credible interval (green hatched region) as reported in Ref. Somasundaram et al. (2024). Though we are not performing a Bayesian analysis, we see very good agreement between our conformal prediction results and the credible interval over all partial wave channels. Interestingly, we do see some slight widening of the uncertainty bands for larger energies in certain channels.
@@ III.3 Phase Shifts
In the previous two subsections, we have considered a single interaction from chiral EFT in order to show the utility of our conformal-prediction approach in the analysis of nucleon-nucleon scattering observables. Here, we instead choose to consider an entire family of local next-to-next-to-leading order (N2LO) interactions Gezerlis et al. (2013, 2014); Somasundaram et al. (2024) that are used in nuclear many-body methods such as quantum Monte Carlo. In Ref. Somasundaram et al. (2024) they report on a Bayesian inference procedure used to fit the nucleon-nucleon interaction and produce posterior distributions for the low-energy couplings (see below). These posterior distributions represent a family of nuclear interactions that all reproduce nucleon-nucleon scattering data. In what follows, we apply our conformal-prediction architecture to nucleon-nucleon phase shifts calculated using this family of interactions. We have also repeated the process for a family of interactions generated ad hoc, where we have assumed naive Gaussian distributions for the low-energy couplings in order to explore how our approach handles different types of parameter distributions. The posterior distributions from Ref. Somasundaram et al. (2024) are compared with our naive Gaussian distributions in Fig. 14, where it can be seen that we have chosen our LECs such that they have similar covariance to the full Bayesian posteriors.
@@
Figure 17: Histogram of phase shifts calculated using the LECs drawn from the Bayesian posteriors in Ref. Somasundaram et al. (2024), at $E=200\ \text{MeV}$ for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves, compared against Gaussian probability density functions. The ${}^{1}S_{0}$ partial wave shows approximately Gaussian behavior, while ${}^{3}S_{1}$ has considerable skewness. The empirical coverage for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves is 0.94844 and 0.94870, respectively for the 2500 samples.
[…]
@@
Figure 15 illustrates both 95% and 68% CP intervals for the phase shifts across a range of energies and different partial waves. These prediction bands are constructed based on samples and reflect the uncertainty in the EFT predictions. These intervals provide uncertainty bands that adapt to the spread of samples. Because conformal prediction is distribution-free, the bands remain valid even in the presence of skewed or non-Gaussian distributions, making them rigorous uncertainty quantification tools in nucleon-nucleon scattering. We can see that by using our conformal-prediction approach we are able to construct uncertainty-bands that capture the data regardless of the underlying distribution and maintain the desired level of empirical coverage. It is interesting to note that the CP bands for the $p$-wave channels seem to be largely unchanged between the two sets of interactions considered, and that the CP bands in the $s$-wave channels show a clear dependence, with considerable broadening for the less-physical family of interactions.
@@
In Fig. 16 we also compare the uncertainty bands generated by our conformal prediction approach against the credible interval results reported in Ref. Somasundaram et al. (2024). While the conformal prediction approach is indeed a post-processing technique, in this case applied to the phase shifts calculated using the posterior distributions from Ref. Somasundaram et al. (2024), it is very reassuring to see that we produce uncertainty bands that agree with these previously published results. We do see some broadening in the larger energies for certain partial waves, which is indicative of how the uncertainty bands change with guaranteed coverage. It would be interesting in future work to attempt a wedding of these two techniques, where perhaps the constraint of imposing a set level of empirical coverage could be imposed in the Bayesian framework.
@@
To further investigate the statistical characteristics of the chiral EFT phase-shift predictions, we analyze the behavior of phase-shift samples found using the Bayesian posterior LECs at E $=200\ \text{MeV}$, for the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves. We visualize their distributions using histograms and compare with Gaussian probability density functions (see Fig. 17). While the ${}^{1}S_{0}$ partial wave appears to have a distribution close to Gaussian, the ${}^{3}S_{1}$ shows skewness or heavy tails. We find similar skewed behavior for both the ${}^{3}D_{1}$ and ${}^{3}P_{2}$ partial waves, indicating that this behavior could be due to the coupled nature of these partial wave channels.
@@
where $\Phi^{-1}$ denotes the quantile function of the standard normal distribution, $n$ is the sample size, and $i$ is the rank of the ordered samples. These values are dimensionless and represent the expected quintiles if the data were normally distributed. The y-axis shows the empirical quantiles of the data, which are simply the ordered phase shift samples at a fixed energy. These values retain their physical units (degrees), as they correspond directly to the sampled phase shift values. The red line in each plot serves as a reference for perfect agreement with the normal distribution. It is typically a straight line fitted to data. If the data are normally distributed, the plotted points lie approximately along the $45$ degree line. Deviations from this line indicate departures from normality. As we can see in Fig. 18, for the ${}^{3}S_{1}$ partial wave, the quantiles deviate significantly from the straight line, particularly in the tails. This deviation shows the presence of skewness in the data. In spite of this skewness, the empirical coverage for the 95% prediction interval at $E=200\ \text{MeV}$ is essentially identical for both the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ partial waves (0.94844 and 0.94870 respectively).
It is important to reiterate that the CP method does not rely on any assumptions about the underlying distribution. CP constructs prediction intervals regardless of whether the data follow a Gaussian or any other distribution.
@@ IV Summary and Outlook
In this work, we developed a distribution-free approach for uncertainty quantification in the nucleon-nucleon scattering problem for physical observables, such as phase shifts and cross sections, using conformal prediction. Unlike Bayesian methods that rely on model assumptions and prior distribution, CP provides finite-sample prediction intervals with guaranteed coverage, regardless of the underlying data distribution. We demonstrated this capability by applying CP to three different scenarios; the pointwise model, the GP model, and phase shifts calculated for local interactions with normally distributed LECs. In the pointwise model, we assumed that the expansion coefficients in the EFT series are sampled from a Student-t distribution, and generated samples for total cross sections at N2LO order at four different energies, then we showed the 95%, and 68% CP bands for the observable. We also calculated the empirical coverage for a range of $1-\alpha$, and showed that our method guarantees the desired coverage. Then, using a Gaussian process, we modeled the expansion coefficients as random functions. The resulting samples at N2LO demonstrated approximately Gaussian behavior, and CP was applied to construct 95% and prediction intervals across different ranges of energies. And finally, in the phase-shift analysis, we explored the behavior of partial- wave samples across different lab energies. Histograms and Q-Q plots revealed that while the ${}^{1}S_{0}$ partial wave followed near Gaussian distribution, the ${}^{3}S_{1}$ shows skewness. Despite these deviations from normality, CP intervals remained valid. By presenting both 68% and 95% CP intervals, we showed how the method can quantify uncertainties. Overall, our results highlight the strength of conformal prediction as a powerful tool for uncertainty quantification. By combining CP with posterior samples, we have established a robust framework with minimal assumptions. We have also checked for both scattering data and nucleon-nucleon phase shifts that our conformal prediction approach provides uncertainty bands that agree well with previously published Bayesian analyses. We hope in the future to build upon this work and develop a framework which marries these two approaches.
@@ Acknowledgments
The authors would like to thank Ingo Tews and Rahul Somasundaram for sharing their LEC posterior distributions and credible interval bands from Ref. Somasundaram et al. (2024), as well as the BUQEYE collaboration for making their source code publicly available. This work was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the Canada Foundation for Innovation (CFI). Computational resources have been provided by Compute Ontario through the Digital Research Alliance of Canada, and by the National Energy Research Scientific Computing Center (NERSC), which is supported by the U.S. Department of Energy, Office of Science, under contract No. DE-AC02-05CH11231.

### v2

@@ III.1 Pointwise Model
To evaluate the validity of the conformal prediction intervals, we plot the empirical coverage to the target coverage levels across a range of $1-\alpha$. The results of this evaluation for the total cross section at E=50MeVE=50\ $\text${MeV} based on $n_{\text{split}}=4000$, are shown in Fig. 12. The empirical coverage values closely follow the ideal line, always remaining within the 68% and 95% confidence band. These results validate the robustness and reliability of our method and show that conformal prediction provides guaranteed coverage with minimal assumptions.
@@ III.2 Gaussian Processes
Here, $l$ is the length scale, which controls how quickly the function can change. In addition, a scaled inverse-chi-squared prior is considered for the $\overline{c}^{2}$, which is a conjugate prior for the variance of the GP. For more details about choosing the priors and hyperparameters in GP, see Refs. Furnstahl et al. (2015b); Melendez et al. (2017, 2019).
Figure 13: CQR intervals for the differential cross section as a function of polar scattering angle, computed at N2LO and E=96MeVE=96\ $\text${MeV} using EKM potential with $R_{0}=0.9\ \text{fm}$. The dark and light blue bands represent the $68\%$ and $95\%$ CQR intervals, respectively. The black curve corresponds to the Nijmegen PWA result Stoks et al. (1993). The prediction bands are obtained by sampling the chiral EFT expansion coefficients using the GP model and applying CQR to construct distribution-free prediction intervals constructed from 10000 samples.
After sampling the total cross-section expansion coefficients at N2LO from the GP model across different ranges of energies, we draw samples for the total cross-section, $\sigma_{tot}$, at each energy value. Using these samples, we then apply CQR to construct a distribution-free prediction interval. In the right panel of Fig. 9 the resulting 68% and 95% CQR bands are shown for the total cross section at N2LO order using the EKM potential with $R_{0}=0.9\ \text{fm}$. The conformal bands widen gradually with increasing energy, capturing the uncertainty associated with the EFT predictions at higher energy. Looking at Fig. 9, we can compare the CQR intervals for the total cross section $\sigma_{tot}$ at N2LO obtained using the pointwise model (left panel) and GP model (right panel), across a range of energies. Both models yield prediction intervals that widen with energy. However, a notable difference is that the GP-based intervals are generally narrower than those from the pointwise model, particularly at higher energies. Additionally, CQR bands based on GP are larger than those of the pointwise model for lower energies. Finally, we note that the bands appear jagged in both cases due to fluctuations in the underlying samples; this effect is stronger for the pointwise model, where samples introduce more noise, and weaker for the GP model, where correlations across energies are taken into account.
To further explore the versatility of CP, we apply it to the differential cross section $\frac{d\sigma}{d\Omega}$ as a function of polar scattering angle at a fixed energy of E=96MeVE=96\ $\text${MeV}. For this, we model the differential cross-section expansion coefficients as random functions using a GP, allowing us to generate samples of the differential cross sections across the different polar angle ranges. The resulting $68\%$ and $95\%$ CQR intervals, shown in Fig. 13, highlight the ability of CP to provide robust, distribution-free uncertainty quantification for a variety of nuclear observables beyond total cross sections. We emphasize that this approach can be generalized to other observables of interest.
@@
Figure 15: 95% and 68% CQR intervals for scattering phase shifts in different partial wave channels as a function of energy, generated using the N2LO interactions as described in the main text. The CQR bands are constructed from 2500 samples and adapt to the spread of the data, with the black curve showing the Nijmegen partial-wave analysis Stoks et al. (1993). Since the underlying raw phase shift samples are smooth, the resulting CQR bands also appear smooth, in contrast to jagged bands observed for the total cross section.
Figure 16: Comparison of our 95% conformal prediction interval (blue shaded region) with the 95% credible interval (green hatched region) as reported in Ref. Somasundaram et al. (2024). We see very good agreement between our conformal prediction results and the credible interval over all partial wave channels. Interestingly, we do see some slight widening of the uncertainty bands for larger energies in certain channels.
@@ III.3 Phase Shifts
In the previous two subsections, we have considered a single interaction from chiral EFT in order to show the utility of our conformal-prediction approach in the analysis of nucleon-nucleon scattering observables. Here, we instead choose to consider an entire family of local next-to-next-to-leading order (N2LO) interactions Gezerlis et al. (2013, 2014); Somasundaram et al. (2024) that are used in nuclear many-body methods such as quantum Monte Carlo. In Ref. Somasundaram et al. (2024) they report on a Bayesian inference procedure used to fit the nucleon-nucleon interaction and produce posterior distributions for the low-energy couplings (see below). These posterior distributions represent a family of nuclear interactions that all reproduce nucleon-nucleon scattering data. In what follows, we apply our conformal-prediction architecture to nucleon-nucleon phase shifts calculated using this group of interactions. We have also repeated the process for a set of interactions generated ad hoc, where we have assumed naive Gaussian distributions for the low-energy couplings in order to explore how our approach handles different types of parameter distributions. The posterior distributions from Ref. Somasundaram et al. (2024) are compared with our naive Gaussian distributions in Fig. 14, where it can be seen that we have chosen our LECs such that they have similar covariance to the full Bayesian posteriors.
@@
Figure 17: Histogram of phase shifts calculated using the LECs drawn from the Bayesian posteriors in Ref. Somasundaram et al. (2024), at E=200MeVE=200\ $\text${MeV} for the $^{1}S_{0}$ and $^{3}S_{1}$ partial waves, compared against Gaussian probability density functions. The $^{1}S_{0}$ partial wave shows approximately Gaussian behavior, while $^{3}S_{1}$ has considerable skewness. The empirical coverage for the $^{1}S_{0}$ and $^{3}S_{1}$ partial waves is 0.94844 and 0.94870, respectively for the 2500 samples.
In Ref. Somasundaram et al. (2024) the LECs at N2LO are fit to reproduce the phase-shift data from the Nijmegen partial-wave analysis (PWA) Stoks et al. (1993) for laboratory energies up to $150$ MeV in the $^{1}\text{S}_{0}$, $^{3}\text{S}_{1}$, $^{1}\text{P}_{1}$, $^{3}\text{P}_{0}$, $^{3}\text{P}_{1}$, and $^{3}\text{P}_{2}$ partial-wave channels and the $\epsilon_{1}$ mixing angle. Therefore, we compute the phase shifts for each of these channels, as well as the coupled 3D1 channel, up to 350MeV350\ $\text${MeV} and use the resulting distribution of phase shifts in order to construct conformal-prediction uncertainty bands.
@@
In Fig. 15 we show both 95% and 68% CQR intervals for the phase shifts across a range of energies and different partial waves. These prediction bands are constructed based on samples and reflect the uncertainty in the EFT predictions. Because conformal prediction is distribution-free, the bands remain valid even in the presence of skewed or non-Gaussian distributions, making them rigorous uncertainty quantification tools in nucleon-nucleon scattering. We can see that by using our conformal-prediction approach we are able to construct uncertainty-bands that capture the data regardless of the underlying distribution and maintain the desired level of empirical coverage. It is interesting to note that the CQR bands for the $p$-wave channels seem to be largely unchanged between the two sets of interactions considered, and that the CQR bands in the $s$-wave channels show a clear dependence, with considerable broadening for the less-physical family of interactions. Finally, it is worth noting that the phase shift intervals show a smooth behavior, which directly reflects the smoothness of the underlying samples. This contrasts with the total cross section case, where noisier samples lead to jagged intervals.
@@
In Fig. 16 we also compare the uncertainty bands generated by our conformal prediction approach against the credible interval results reported in Ref. Somasundaram et al. (2024). While the conformal prediction approach is indeed a post-processing technique, in this case applied to the phase shifts calculated using the posterior distributions from Ref. Somasundaram et al. (2024), it is very reassuring to see that we produce uncertainty bands that agree with these previous results. We do see some broadening in the larger energies for certain partial waves, which is indicative of how the uncertainty bands change with guaranteed coverage. It would be interesting in future work to attempt a wedding of these two techniques, where perhaps the constraint of imposing a set level of empirical coverage could be imposed in the Bayesian framework.
@@
To further investigate the statistical characteristics of the chiral EFT phase-shift predictions, we analyze the behavior of phase-shift samples found using the Bayesian posterior LECs at E =200MeV=200\ $\text${MeV}, for the $^{1}S_{0}$ and $^{3}S_{1}$ partial waves. We visualize their distributions using histograms and compare with Gaussian probability density functions (see Fig. 17). While the $^{1}S_{0}$ partial wave appears to have a distribution close to Gaussian, the $^{3}S_{1}$ shows skewness or heavy tails. We find similar skewed behavior for both the $^{3}D_{1}$ and $^{3}P_{2}$ partial waves, indicating that this behavior could be due to the coupled nature of these partial wave channels.
@@
where $\Phi^{-1}$ denotes the quantile function of the standard normal distribution, $n$ is the sample size, and $i$ is the rank of the ordered samples. These values are dimensionless and represent the expected quintiles if the data were normally distributed. The y-axis shows the empirical quantiles of the data, which are simply the ordered phase shift samples at a fixed energy. These values retain their physical units (degrees), as they correspond directly to the sampled phase shift values. The red line in each plot serves as a reference for perfect agreement with the normal distribution. It is typically a straight line fitted to data. If the data are normally distributed, the plotted points lie approximately along the $45$ degree line. Deviations from this line indicate departures from normality. As we can see in Fig. 18, for the $^{3}S_{1}$ partial wave, the quantiles deviate significantly from the straight line, particularly in the tails. This deviation shows the presence of skewness in the data. In spite of this skewness, the empirical coverage for the 95% prediction interval at E=200MeVE=200\ $\text${MeV} is essentially identical for both the $^{1}S_{0}$ and $^{3}S_{1}$ partial waves (0.94844 and 0.94870 respectively). It is important to reiterate that the CP method does not rely on any assumptions about the underlying distribution. CP constructs prediction intervals regardless of whether the data follow a Gaussian or any other distribution.
@@ IV Summary and Outlook
In this work, we developed a distribution-free approach for uncertainty quantification in the nucleon-nucleon scattering problem for physical observables, such as phase shifts and cross sections, using conformal prediction. Unlike Bayesian methods that rely on model assumptions and prior distribution, CP provides finite-sample prediction intervals with guaranteed coverage, regardless of the underlying data distribution. We demonstrated this capability by applying CP to three different scenarios; the pointwise model, the GP model, and phase shifts calculated for local interactions with normally distributed LECs. In the pointwise model, we assumed that the expansion coefficients in the EFT series are sampled from a Student-t distribution, and generated samples for total cross sections at N2LO order at four different energies, then we showed the 95% and 68% CQR bands for the observable. We also calculated the empirical coverage for a range of $1-\alpha$, and showed that our method guarantees the desired coverage. Then, using a Gaussian process, we modeled the expansion coefficients as random functions. The resulting samples at N2LO demonstrated approximately Gaussian behavior, and CQR was applied to construct 95% and prediction intervals across different ranges of energies. And finally, in the phase-shift analysis, we explored the behavior of partial- wave samples across different lab energies. Histograms and Q-Q plots revealed that while the $^{1}S_{0}$ partial wave followed near Gaussian distribution, the $^{3}S_{1}$ shows skewness. Despite these deviations from normality, CQR intervals remained valid. By presenting both 68% and 95% CQR intervals, we showed how the method can quantify uncertainties. Overall, our results highlight the strength of conformal prediction as a powerful tool for uncertainty quantification. By combining CP with posterior samples, we have established a robust framework with minimal assumptions. We have also checked for both scattering data and nucleon-nucleon phase shifts that our conformal prediction approach provides uncertainty bands that agree well with previously published Bayesian analyses. We hope in the future to build upon this work and develop a framework which marries these two approaches.
@@ Acknowledgments
The authors would like to thank Ingo Tews and Rahul Somasundaram for sharing their LEC posterior distributions and credible interval bands from Ref. Somasundaram et al. (2024), as well as the BUQEYE collaboration for making their source code publicly available. We would also like to thank ECT* for support at the Workshop Next generation ab initio nuclear theory during which this work was discussed. This work was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the Canada Foundation for Innovation (CFI). Computational resources have been provided by Compute Ontario through the Digital Research Alliance of Canada, and by the National Energy Research Scientific Computing Center (NERSC), which is supported by the U.S. Department of Energy, Office of Science, under contract No. DE-AC02-05CH11231.